OpenStudy (anonymous):

Christa will drive sixty-five miles per hour for the next several hours. If she is already ninety miles away from her house, how many miles away from her house will she be after two hours? The expression is ninety plus sixty-five times h, where h is the number of hours she drives. one hundred thirty two hundred twenty two hundred forty-five two hundred sixty

OpenStudy (anonymous):

90 + 2*65 = 90 + 130 = 220

OpenStudy (anonymous):

thankz